OpenStudy (anonymous):

The function f is defined as follows f(x) { 3+x if x< 0 {x^2 if x > 0 (a) Find the domain of the function (b) Locate any intercepts (c) Graph the functions (d) Based on the graph, find the range (e) Is f continuous on its domain?

OpenStudy (jdoe0001):

ok, so is a "piecewise" function, it has 2 patterns so let's find the domain for f(x) { 3+x if x< 0 {\(x^2\) if x > 0 what values can "x" take based on those "if" restrictions?

OpenStudy (anonymous):

I have no idea is it 0

OpenStudy (jdoe0001):

well, can we give x = 25?

OpenStudy (jdoe0001):

is there anything there saying we can't?

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

or say let's make x = 15, or 18, or -25, or -15, or -18 is there anything saying we can't?

OpenStudy (anonymous):

no ma'am

OpenStudy (jdoe0001):

the function f(x) { 3+x if x< 0 # x is less than 0, not equal to 0, just less, how less? up to \(-\infty\) it seems {x2 if x > 0 # x is greater than 0, not equal to 0, just greater, how much bigger? up to \(+\infty\) it seems

OpenStudy (jdoe0001):

so, x can go to infinity on either side, BUT never be 0 so, that's the DOMAIN

OpenStudy (jdoe0001):

or in interval notation will be \((-\infty, 0)\) and \((0, +\infty)\)

OpenStudy (jdoe0001):

open intervals, because, it never gets to 0, and infinity is really never reachable either

OpenStudy (anonymous):

okay @ jdoe, the answer is (00,0) and (0,00)? I know what the double o's, cant implement it on the keyboard

OpenStudy (jdoe0001):

see above :), yes

OpenStudy (anonymous):

The domain can be expressed as {x | x < 0} u {x|x>0} too

OpenStudy (anonymous):

okay thanks guys!!!

OpenStudy (jdoe0001):

so, since x will never become 0, that means, you won't have a y-intercept, because y-intercepts occur only when x = 0, and in this case, x cannot be =0 what about x intercepts? x intercepts occur when y is set to 0 so let's check the 1st one y = 3+x 0 = 3+x => x = -3 # so that's one x intercept, at (-3, 0) let's check the 2nd one y = \(x^2\) 0 = \(x^2 \implies x =0\) # no dice because x cannot be =0, we're restricted there # so no x intercepts on the 2nd one, just the 1st one

OpenStudy (anonymous):

okay.... got it

OpenStudy (jdoe0001):

now you'd just need to graph them, keeping in mind their range so it looks more or less like 

OpenStudy (anonymous):

Wow thanks, you learn something new everyday...:)

OpenStudy (jdoe0001):

is a "piecewise" function, so it has 2 patterns, one for x<0, and another pattern for x>0 is really an inclined line when x<0 and is a parabola when x> 0

OpenStudy (anonymous):

okay....

OpenStudy (jdoe0001):

based on the graph, let's find the range what values is "y" taking? the inclined line on the left is going and going down and down further down to \(-\infty\) the parabola on the right, is going up and up and up and up further up to \(+\infty\)

OpenStudy (jdoe0001):

and that's the range, "y" is really "occupying" from \((-\infty, +\infty)\)

OpenStudy (jdoe0001):

now, if the piecewise function "continuous" as in in a single continum, as in an unbroken thread well, just look at the graph, that'd tell you :)

OpenStudy (anonymous):

okay thanks jdoe your the best!!!

OpenStudy (jdoe0001):

yw

OpenStudy (whpalmer4):

@charisse1 if you ever see something that you don't know how to make, select it, then right-click and choose Show Math As>TeX Commands and you'll get a little pop-up menu that tells you what you need to type in LaTeX to generate it. For example, \[(-\infty, +\infty)\]is created with (-\infty, +\infty)

OpenStudy (whpalmer4):

uh, not a pop-up menu, a pop-up window...